OpenStudy (anonymous):

Jackson cements 24 triangular tiles behind his sink. The base and height of each tile is about 6 inches. How much area in feet does he cover?

hartnn (hartnn):

area of 1 tile =0.5* base * height = 0.5*6*6 so area of 24 tiles= 24*0.5*6*6 = ??

OpenStudy (anonymous):

i got 432

hartnn (hartnn):

the answer u will get in inch^2 just convert it in feet^2

OpenStudy (anonymous):

so it'll be 432ft^2

hartnn (hartnn):

no u need to convert 432 inch^2=??? ft^2

OpenStudy (anonymous):

am i multiplying 432 x 2?

hartnn (hartnn):

nopes, just divide it by (12*12)

hartnn (hartnn):

432/(12*12)=432/144=?

OpenStudy (anonymous):

i got 3ft^2

hartnn (hartnn):

because 1 foot is 12 inch thats correct 3 ft^2